OpenStudy (anonymous):

Find a cubic function with the given zeros.√(2), -√(2), -2

OpenStudy (anonymous):

\[(x-\sqrt{2})(x-(-\sqrt{2}))(x-(-2))\]

OpenStudy (anonymous):

and multiply

OpenStudy (anonymous):

i did but the answer didnt come up and i have these too pick from

OpenStudy (anonymous):

f(x) = x3 + 2x2 - 2x + 4 f(x) = x3 + 2x2 + 2x - 4 f(x) = x3 - 2x2 - 2x - 4 f(x) = x3 + 2x2 - 2x - 4

OpenStudy (anonymous):

it is one of them. try again